# Spatio-Temporal Fuzzy-oriented Multi-Modal Meta-Learning for Fine-grained Emotion Recognition

### v2

@@ V-E Ablation Study
V-F Parameter Investigation
The impact of hyperparameters on performance cannot be overlooked. In the case of ST-F2M, the description of membership functions in FCIS and FKIS directly affects emotion recognition results. Therefore, we conduct experiments specifically focusing on the attribute membership range $\lambda_{1}$ in FCIS and the membership range $\lambda_{2}$ in FKIS to find the optimal settings. The example results are presented in a 3D histogram in Figure 10. Example results are presented as a 3D histogram in figure 10. Specifically, we take the high-intensity value sample of angry emotion as an example to calculate the membership range of the generalized fuzzy inference module when the accuracy and recall rates are the highest. The best effect is when $\lambda_{1}$ and $\lambda_{2}$ are both 0.4, which is also the setting of our proposed fuzzy systems. The results indicate that the membership functions we constructed for the system in Subsection III-B deliver the best performance. This demonstrates the rigor of our work.
V-G Practical Case Study
@@
Experimental Setup
We use the same implementation details as Subsection V-A but a different deployment platform, i.e., a self-built robot platform  [^1] More details on the robot build, please see: https://github.com/WangJingyao07/PickBallBot-5mao/tree/main. The robot host we use is NVIDIA Jetson AGX Xavier, including a 512-core NVIDIA Volta GPU with 64 Tensor cores, two NVIDIA deep learning accelerator pedals, two visual accelerator pedals, and an eight-core NVIDIA Carmel Arm CPU. Our ST-F2M is experimented on a unified model edge computing platform and ultimately embedded on a medical robotics platform. We set up the communication and underlying control algorithms to achieve a complete system-closed loop. In this experiment, we use the built robot entity to conduct 5 groups of experiments in the laboratory simulation environment. The data of each group contains 20 pieces of data with different angles. Figure 11(a) shows the data collection. Each piece of data is a 30-second video, labeled with the emotional sequence contained in the video. The robot’s goal is to recognize the full range of emotions contained in video data. Note that the data in this scene may have various interferences, such as camera distortion and object occlusion [100, 101]. Figure 1(c) provides examples of FER samples in real-life scenarios.
Results
Figure 11 provides the experimental results. From the results, we can observe that: (i) ST-F2M still achieves superior results in real-life deployment; (ii) other FER methods have unsatisfactory results in the robot entity test environment, possibly due to various reasons like camera distortion, object occlusion, etc.; and (iii) ST-F2M can still obtain relatively stable FER results when facing many problems caused by the robot entity testing environment. These observations prove the advantages of ST-F2M in practical applications and illustrate the effectiveness and forward-lookingness of our design for the real-life FER field.
VI Conclusion

### v4

@@ VI-E Ablation Study
VI-F Parameter Investigation
The impact of hyperparameters on performance cannot be overlooked. In the case of ST-F2M, the description of membership functions in FCIS and FKIS directly affects emotion recognition results. Therefore, we conduct experiments specifically focusing on the attribute membership range $\lambda_{1}$ in FCIS and the membership range $\lambda_{2}$ in FKIS to find the optimal settings. The example results are presented in a 3D histogram in Figure 10. Example results are presented as a 3D histogram in figure 10. Specifically, we take the high-intensity value sample of angry emotion as an example to calculate the membership range of the generalized fuzzy inference module when the accuracy and recall rates are the highest. The best effect is when $\lambda_{1}$ and $\lambda_{2}$ are both 0.4, which is also the setting of our proposed fuzzy systems. The results indicate that the membership functions we constructed for the system in Subsection IV deliver the best performance. This demonstrates the rigor of our work.
VI-G Practical Case Study
@@
We use the same implementation details as Subsection VI-A but a different deployment platform, i.e., a self-built robot platform  [^1] More details on the robot build, please see: https://github.com/WangJingyao07/PickBallBot-5mao/tree/main. The robot host we use is NVIDIA Jetson AGX Xavier, including a 512-core NVIDIA Volta GPU with 64 Tensor cores, two NVIDIA deep learning accelerator pedals, two visual accelerator pedals, and an eight-core NVIDIA Carmel Arm CPU. Our ST-F2M is experimented on a unified model edge computing platform and ultimately embedded on a medical robotics platform. We set up the communication and underlying control algorithms to achieve a complete system-closed loop. In this experiment, we use the built robot entity to conduct 5 groups of experiments in the laboratory simulation environment. The data of each group contains 20 pieces of data with different angles. Figure 11(a) shows the data collection. Each piece of data is a 30-second video, labeled with the emotional sequence contained in the video. The robot’s goal is to recognize the full range of emotions contained in video data. Note that the data in this scene may have various interferences, such as camera distortion and object occlusion [100, 101]. Figure 1(c) provides examples of real-world FER samples.
From Figure 11, we can observe that: (i) ST-F2M still achieves superior results in real-life deployment; (ii) other FER methods have unsatisfactory results in the robot entity test environment, possibly due to various reasons like camera distortion, object occlusion, etc.; and (iii) ST-F2M can still obtain relatively stable FER results when facing many problems caused by the robot entity testing environment. These observations prove the advantages of ST-F2M in practical applications and illustrate the effectiveness and forward-lookingness of our design for the real-life FER field.
VII Conclusion